If Ram walks at 5 kmph to his school then he reaches 5 minutes early and if he walks at 4 kmph he reaches 5 minutes late . what is the actual distance Ram's house and his school?
Answer by mananth(16946)

If Ram walks at 5 kmph to his school then he reaches 5 minutes early and if he walks at 4 kmph he reaches 5 minutes late . what is the actual distance Ram's house and his school?
let distance be x
Let his original speed be y to reach in time
5 minutes = 5/60 hours = 1/12 hours
Normal time = x/y
x/y -x/5 = 1/12
x/4 -x/y = 1/12
add the two equations
x/4 -x/5 = 1/6
x/20 = 1/6
6x=20
x= 10/3
distance = 3.33 km